A quadratic equation has an x intercept at 2.5 and 5 and a y value y=-2.25. Use any valid method to calulate the equation of the quadratic in the form y=ax^2 +bx+c
Answer by lwsshak3(11628)

A quadratic equation has an x intercept at 2.5 and 5 and a y value y=-2.25. Use any valid method to calulate the equation of the quadratic in the form y=ax^2 +bx+c
***
(x-2.5)(x-5)=-2.25
x^2-7.25x+12.5=-2.25
x^2-7.25x+14.75=0